OpenStudy (anonymous):

Suppose p varies directly with d, and p = 7 when d = 3. What is the value of d when p = 28? A. 140/3 B. 4/3 C. 16 D. 12

OpenStudy (anonymous):

ill give you a metal!

OpenStudy (anonymous):

\[\frac{ 7 }{ 3 }= \frac{ 28 }{x}\] 28•3/7=12

OpenStudy (anonymous):

thanks!